# Evaluating Conversational Recommender Systems via Large Language Models

### v3

@@ Abstract.
In this work, we propose a user-centric evaluation framework based on LLMs for CRS, namely Conversational Recommendation Evaluator (CoRE [^1] This article is a follow-up study of Chen et al. (2025), which has been accepted by HCRS@TheWebConf 2025 but has not been publicly published. Sections 4.1 and 7.2 of this paper are based on Chen et al. (2025), and the remaining sections comprise newly introduced research content.. CoRE consists of two main components: (1) LLM-As-Evaluator. Firstly, we comprehensively summarize 12 key factors influencing user experience in CRSs and directly leverage LLM as evaluator to assign a score to each factor. (2) Multi-Agent Debater. Secondly, we design a multi-agent debate framework with four distinct roles (common user, domain expert, linguist, and HCI expert) to discuss and synthesize the 12 evaluation factors into a unified overall performance score.
@@
ccs: Information systems Evaluation of retrieval resultsccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Figure 1. The overview of our method. We evaluate user interactions with conversational systems through a large language model across 12 factors. The scores and justifications for the 12 factors are used as input to the Multi-Agent Debater, where each agent receives inputs corresponding to the specific factors associated with its designated role. After several rounds of debate, the agents produce a final overall score.
@@
Unlike traditional Recommender Systems (RSs), which model user interests based on implicit feedback such as click-through rates or browsing history and generate item recommendations accordingly (Schafer et al., 2007; Ekstrand et al., 2011; Koren, 2008), Conversational Recommender Systems (CRSs) employ language models to collect and interpret explicit user feedback through dialogue interactive dialogue to collect explicit user feedback. Notably, CRSs not only generate natural language responses but can also present a separate list of recommended items (Jannach et al., 2021; Gao et al., 2021). Figure 2 illustrates a multi-turn interaction where the user requests horror movie recommendations and information about a director, and the system responds accordingly. Based on the dialogue, a separated list of inferred relevant items is displayed on the right.
@@
The intrinsic complexity of the conversational recommendation task, which stems from the integration of dialogue management and recommendation generation, not only distinguishes CRSs from both traditional recommender systems and conventional dialogue systems, but also presents substantial challenges for the evaluation of CRSs (Jannach, 2022). On the one hand, CRSs must purposefully drive users to express their potential needs in dialogue, ensuring the natural and comfortable user dialogue experience; and on the other hand, CRSs also need to infer the user’s item preferences through dialogue and provide accurate recommendations in conversation and in the separated list. Nevertheless, current evaluation practices of CRSs fail to address the integrated nature of dialogue flows manegemtn and user preferences modeling, instead treating item recommendation and dialogue management as isolated tasks. Specifically, most studies adopt task-specific rule-based metrics, such as Recall for recommendation accuracy and BLEU for dialogue quality  (Chen et al., 2019; Wang et al., 2022a, b; Zhang et al., 2024b; Feng et al., 2023). This ignores the nature of conversational recommendation in which dialogue management is integrated with user preference modelling, resulting in a biased evaluation.
The first issue of most of the current evaluation paradigm is fragmented evaluation.The fragmented evaluation approach overlooks the interdependencies between recommendation and conversational dynamics, limiting its ability to holistically and fairly evaluate CRSs performance. This shortcoming directly hinders researchers from fairly evaluating the system’s overall performance. The fragemented evaluation methods truggle to address the practical overall CRSs evaluation problems in reality:
@@
Another issue of most of the current evaluation paradigm is misalignment with user experience. Empirical studies argued that rule-based evaluation metrics, such as Recall and BLEU, often fail to align with actual user experiences (Reiter, 2018; Chen et al., 2017; Cremonesi et al., 2012; Manzoor et al., 2024). These metrics are insufficient for accurately reflecting the quality of conversational recommendations and may instead introduce biased or unreasonable estimates (Rosario and Noever, 2023).
@@
Some prior research in dialogue system and CRS evaluation has identified several key factors, such as accuracy, novelty, proactiveness and explainability, that influence user experience through empirical studies (Cai and Chen, 2020; Cai et al., 2022; Siro et al., 2023; Jin et al., 2021, 2024); Meanwhile, some studies (e.g., Sakai (2023); Langevin et al. ([n. d.])) has proposed abstract frameworks aiming to integrate multiple dimensions of system performance into a unified assessment that can be used for CRS evaluation. But these efforts have yet to yield standardized metrics for the direct assessment of CRS performance.
@@
On the other hand, recent studies argue that Large Language Models (LLMs), demonstrating significant natural language understanding capabilities, have significant potential to align with human preferences in text quality evaluation (Liu et al., 2023; Fu et al., 2024; Chen et al., 2023b; Chiang and Lee, 2023a; Wang et al., 2024; Gao et al., 2023). Thus, this positions LLMs as a promising tool for intelligent evaluation of CRSs. However, despite this potential, only a few studies (Wang et al., 2023a; Huang et al., 2024) have explored LLM-based evaluation for conversational recommendation tasks. Moreover, none of these studies have addressed the holistic assessment of CRSs integrating recommendation with dialogue management, leaving a gap in methodologies for unified performance measurement.
@@
To address this gap, in this work, we propose a user-centric evaluation framework based on LLMs for CRSs, namely Conversational Recommendation Evaluator (CoRE). As shown in Figure 1, the framework consists of of two sequential components, with the upper right part represents the first part, while the lower section corresponds to the second part.
@@
The first part is LLM-As-Evaluator: LLMs are employed as evaluators to assess conversation logs between users and CRSs. The evaluation focuses on 12 key factors influencing user experience in conversational recommendation, such as coherence and explainability. Each factor is assigned a score ranging from 0 to 4, accompanied by detailed reasoning. This component is depicted in the top right part of Figure 1.
@@
The second part is Multi-Agent Debater: We design a multi-agent debate framework to derive a unified overall performance score based on the evaluation scores and explanations of the 12 factors. Specifically, four agents with distinct roles (i.e., a common user, a domain expert, a linguist, and an HCI expert) representing evaluators with different background and specialized knowledge, engage in discussions and collaboratively determine a final overall score ranging from 0 to 100. This design overcomes the limitations of relying solely on a single annotator, which may introduce bias and instability, and aligns with best practices that emphasize collaboration among multiple annotators (van der Lee et al., 2019; Karpinska et al., 2021a). This component is depicted in the bottom part of Figure 1.
@@
  • RQ2: To what extent are the evaluation results of the 12 key factors by LLMs consistent with those provided by human evaluators?
@@
To answer the above research questions, we conducted comprehensive experiments on two widely-used benchmark datasets, ReDial (Li et al., 2019) and OpenDialKG (Moon et al., 2019), involving four representative CRSs , including both neural-based models (BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), UniCRS (Wang et al., 2022b)) and a GPT-based model (CHATCRS (Wang et al., 2023a)). We adopted a modified user simulator based on Wang et al. (2023a) to generate multi-turn dialogue logs that realistically mimicked user interactions by providing vague requests and ensuring a dynamic conversation. The generated logs were then evaluated using three large language models (GPT-4o-mini, LLaMa-3-8B, and GLM-4-Air) across the twelve human-centered factors. Then the Multi-Agent Debater based on GPT-4o-mini was introduced to aggregates factor-wise scores provided by GPT-4o-mini into a unified overall assessment score. Additionally, to validate the proposed framework, we conducted a human evaluation where ten independent annotators rated a total of 80 randomly selected dialogue logs from the test sets. Each evaluator scored the logs on the same twelve factors and provided an overall score, allowing for direct comparison with the LLM-based assessments.
From experimental results, we have the following key findings: (1) The key factors in the proposed framework show a degree of independence under human evaluation, but LLM based evaluators tend to blur those boundaries to different degrees: some models introduce only moderate coupling, while others create much stronger inter factor interdependencies. (2) The proposed framework (CoRE), with appropriated LLM serving as the evaluator, aligns well with human evaluation in most of the key factors influencing user experience in conversational recommendations. (3) CoRE, with appropriated LLM serving as the evaluator, aligns well with human evaluation in terms of CRS overall performance evaluation. (4) LLMs have strong potential of being used directly as CRS. Neural network-based CRS could exhibit deficiencies in some key user experience factors such as Semantic Relevance, Explainability, and Proactiveness, which negatively impact the overall user interaction experience.
@@ 2.1. Conversational Recommender Systems
Dialogue systems are designed to provide social support or functional services to human users through natural language interactions. Key areas in this field has primarily focused on dialogue context comprehension (Chen et al., 2022a) and response generation techniques (Roller et al., 2021). Recommender Systems aim to help users filter and select items that meet their information needs. To achieve this goal, traditional recommender systems generate recommendations based on implicit feedbacks from past user-item interactions (Schafer et al., 2007; Ekstrand et al., 2011; Koren, 2008; Bell and Koren, 2007; Koren et al., 2009; Covington et al., 2016). As the pace of modern life accelerates, traditional recommender systems face inherent limitations due to their static interaction paradigms and insufficient adaptability to users’ evolving preferences. In response, Conversational Recommender Systems (CRSs) have emerged as a promising research frontier, integrating intelligent dialogue agents with recommendation engines to enable dynamic, context-aware, and user-centric interactions (Ren et al., 2020). Taking a different approach to traditional RSs, CRSs aim to model the relationship between user preferences and candidate item representations to deliver accurate recommendations , while simultaneously generating coherent responses that are contextually aligned with the recommended items (Jannach et al., 2021; Gao et al., 2021; Zamani et al., 2023; Li et al., 2018; Liang et al., 2021). As CRSs interact with users through multi-turn dialogues, CRSs are also regarded as a specialized form of dialogue systems (Deng et al., 2025; Li et al., 2021; Ren et al., 2020).
@@
In our study, we focus on the evaluation of chit-chat based CRS, and we selected four representative CRSs for evaluation, in which three (BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019) and UniCRS (Wang et al., 2022b)) are based on neural networks, while one CHATCRS (Wang et al., 2023a)) directly utilizes GPT as a CRS.
@@
We selected the above systems because many studies (Wang et al., 2023a, b; Zhu et al., 2024b, a) have evaluated them or used them as baselines, which helps to connect our work with existing evaluations of conversational recommender systems.
@@ 2.2. Evaluating Conversational Recommender Systems
Evaluation is a fundamental aspect of developing search and recommender systems, encompassing both user-centric assessments, which captures the effectiveness of the system from the user’s perspective, such as user satisfaction, as well as system-centric metrics such as Precision, Recall, and inference time (Siro et al., 2023; Al-Maskari and Sanderson, 2010). The evaluation of CRS also encompasses both user-centric and system-centric approaches, and can be further categorized into three main scenarios:
  • Lab-based User studies, examine perceived system quality, evaluating aspects like recommendation relevance, perceived transparency, and ease of use (Pu et al., 2011).
  • Online metrics, such as A/B testing is commonly used in real-world settings to assess whether the system achieves its intended objectives, typically through metrics such as user engagement or conversion rates (Jannach and Jugovac, 2019).
@@
Due to the high cost of A/B testing and user studies, offline evaluation is commonly adopted in practice to assess CRSs. The most commonly used metrics for evaluating CRS include Recall, Precision, etc., for assessing recommendation task performance, and BLEU, ROUGE, Perplexity, Distinct-N, etc., for evaluating dialogue management task performance. Nevertheless, empirical studies indicate that these rule-based metrics do not align with human preferences (Chen et al., 2017; Reiter, 2018; Cremonesi et al., 2012). One key contribution of this study is the use of large language models (LLMs) as evaluators to enable user-centric offline evaluation.
@@
In the context of user-centered evaluation for CRSs, previous studies have argued that several factors may influence users’ assessment of their interactions with the system, particularly from both the recommendation and dialogue perspectives . These factors include: (perceived) accuracy of recommendation (Contreras et al., 2021; Manzoor et al., 2024; Jin et al., 2021; Zhang et al., 2024a), diversity (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024a) of recommended items, novelty of recommended items (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024a), explainability of recommendation (Sakai, 2023; Jin et al., 2021; Zhang et al., 2024a, c), user intent understanding and response capability (e.g., maintaining a coherent dialogue flow, recovering from errors) (Sakai, 2023; Langevin et al., [n. d.]; Jin et al., 2021; Zhang et al., 2024a), the quality of system-generated text (e.g., fluency, linguistic appropriateness and naturalness) (Jin et al., 2021; Jannach, 2022; Langevin et al., [n. d.]), system informativeness for decision-making (e.g., information sufficiency, information accuracy) (Sakai, 2023; Jannach, 2022), proactiveness of the system (Kraus et al., 2020; Deng et al., 2025; Christakopoulou et al., 2016). Nevertheless, these studies did not propose any computable metric for evaluating CRS performance.
@@
Some prior work has explored both human-annotated and LLM-based approaches for evaluating CRSs from a user-centered perspective.  Manzoor and Jannach (2022) uses crowd-sourced pairwise comparisons to assess CRS response quality, and Yang et al. (2024) proposed a behavior alignment method that relies on human-annotated data. However these methods rely on collecting human data, which is costly.  Wang et al. (2023a) utilized LLMs to assess whether the items recommended by CRS meet user needs, while  Huang et al. (2024) leveraged LLMs to evaluate the social attributes of CRS.  Zhang et al. (2024c) leveraged LLMs to assess the quality of explanation texts generated by CRSs. However, these studies did not proposed a method for assessing the overall performance of CRS. To the best of our knowledge, our work is the first to propose a unified, LLM-based framework for holistic, user-centered evaluation of CRSs. Our work bridges the gap between user studies on CRSs and current evaluation practices in the field. We extracted twelve factors influencing users’ interaction experiences with conversational recommender systems from prior studies, used LLMs to assign evaluation scores to each factor, and then aggregated these into an overall score. By proposing a feasible evaluation framework, our study contributes to advancing CRS evaluation practices toward a more human-centered paradigm.
2.3. Large Langugae Models as Evaluators for Text Quality

### v4

@@ Abstract.
In this work, we propose a user-centric evaluation framework based on LLMs for CRS, namely Conversational Recommendation Evaluator (CoRE [^1] This article is a follow-up study of Chen et al. (2025a), which has been accepted by HCRS@TheWebConf 2025 but has not been publicly published. Sections 4.1 and 7.2 of this paper are based on Chen et al. (2025a), and the remaining sections comprise newly introduced research content.. CoRE consists of two main components: (1) LLM-As-Evaluator. Firstly, we comprehensively summarize 12 key factors influencing user experience in CRSs and directly leverage LLM as evaluator to assign a score to each factor. (2) Multi-Agent Debater. Secondly, we design a multi-agent debate framework with four distinct roles (common user, domain expert, linguist, and HCI expert) to discuss and synthesize the 12 evaluation factors into a unified overall performance score.
@@
$c$s: Information systems Evaluation of retrieval resultsccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Figure 1. The overview of our method. We evaluate user interactions with conversational systems through a large language model across 12 factors. The scores and justifications for the twelve factors are used as input to the Multi-Agent Debater, where each agent receives inputs corresponding to the specific factors associated with its designated role. After several rounds of debate, the agents produce a final overall score.
@@
Unlike traditional Recommender Systems (RSs), which model user interests based on implicit feedback such as click-through rates or browsing history and generate item recommendations accordingly (Schafer et al., 2007; Ekstrand et al., 2011; Koren, 2008), Conversational Recommender Systems (CRSs) employ language models to collect and interpret explicit user feedback through interactive dialogue to collect explicit user feedback. Notably, CRSs not only generate natural language responses but can also present a separate list of recommended items (Jannach et al., 2021; Gao et al., 2021). Figure 2 illustrates a multi-turn interaction where the user requests horror movie recommendations and information about a director, and the system responds accordingly. Based on the dialogue, a separated list of inferred relevant items is displayed on the right.
@@
The intrinsic complexity of the conversational recommendation task, which stems from the integration of dialogue management and recommendation generation, not only distinguishes CRSs from both traditional recommender systems and conventional dialogue systems, but also presents substantial challenges for the evaluation of CRSs (Jannach, 2022). On the one hand, CRSs must purposefully drive users to express their potential needs in dialogue, ensuring the natural and comfortable user dialogue experience; and on the other hand, CRSs also need to infer the user’s item preferences through dialogue and provide accurate recommendations in conversation and in the separated list. Nevertheless, current evaluation practices of CRSs fail to address the integrated nature of dialogue flow management and user preferences modeling, instead treating item recommendation and dialogue management as isolated tasks. Specifically, most studies adopt task-specific rule-based metrics, such as Recall for recommendation accuracy and BLEU for dialogue quality  (Chen et al., 2019; Wang et al., 2022a, b; Zhang et al., 2024a; Feng et al., 2023). This ignores the nature of conversational recommendation in which dialogue management is integrated with user preference modelling, resulting in a biased evaluation.
The first issue with most current evaluation paradigms is fragmented evaluation. The fragmented evaluation approach overlooks the interdependencies between recommendation and conversational dynamics, limiting its ability to holistically and fairly evaluate CRSs’ performance. This shortcoming directly hinders researchers from fairly evaluating the system’s overall performance. Specifically, the fragmented evaluation methods struggle to address the practical overall CRSs evaluation problem:
@@
Another issue with most current evaluation paradigms is their misalignment with user experience. Empirical studies argued that rule-based evaluation metrics, such as Recall and BLEU, often fail to align with actual user experiences (Reiter, 2018; Chen et al., 2017; Cremonesi et al., 2012; Manzoor et al., 2024). These metrics are insufficient for accurately reflecting the quality of conversational recommendations and may instead introduce biased or unreasonable estimates (Rosario and Noever, 2023).
@@
Some prior research in dialogue systems and CRS evaluation has identified several key factors, such as accuracy, novelty, proactiveness and explainability, that influence user experience through empirical studies (Cai and Chen, 2020; Cai et al., 2022; Siro et al., 2023; Jin et al., 2021, 2024). Meanwhile, some studies (e.g., Sakai (2023); Langevin et al. (2021)) have proposed abstract frameworks aiming to integrate multiple dimensions of system performance into a unified assessment that can be used for CRS evaluation. But these efforts have yet to yield standardized metrics for the direct assessment of CRS performance.
@@
On the other hand, recent studies argue that Large Language Models (LLMs), demonstrating considerable natural language understanding capabilities, have significant potential to align with human preferences in text quality evaluation (Liu et al., 2023; Fu et al., 2024; Chen et al., 2023b; Chiang and Lee, 2023a; Wang et al., 2024; Gao et al., 2023). Thus, this positions LLMs as a promising tool for intelligent evaluation of CRSs. However, despite this potential, only a few studies (Wang et al., 2023a; Huang et al., 2024) have explored LLM-based evaluation for conversational recommendation tasks. Moreover, none of these studies have addressed the holistic assessment of CRSs integrating recommendation with dialogue management, leaving a gap in methodologies for unified performance measurement.
@@
To address this gap, in this work, we propose a user-centric evaluation framework based on LLMs for CRSs, namely Conversational Recommendation Evaluator (CoRE). As shown in Figure 1, the framework consists of two sequential components, with the upper right part represents the first part, and the lower section representing the second part.
@@
The first part is LLM-As-Evaluator: LLMs are employed as evaluators to assess conversation logs between users and CRSs. The evaluation focuses on twlve key factors influencing user experience in conversational recommendation, such as coherence and explainability. Each factor is assigned a score ranging from 0 to 4, accompanied by detailed reasoning. This component is depicted in the top right part of Figure 1.
@@
The second part is Multi-Agent Debater: We design a multi-agent debate framework to derive a unified overall performance score based on the evaluation scores and explanations of the twelve factors. Specifically, four agents with distinct roles (i.e., a common user, a domain expert, a linguist, and an HCI expert) representing evaluators with different backgrounds and specialized knowledge, engage in discussions and collaboratively determine a final overall score ranging from 0 to 100. This design overcomes the limitations of relying solely on a single annotator, which may introduce bias and instability, and aligns with best practices that emphasize collaboration among multiple annotators (van der Lee et al., 2019; Karpinska et al., 2021a). This component is depicted in the bottom part of Figure 1.
@@
  • RQ2: To what extent are the evaluation results of the twelve key factors by LLMs consistent with those provided by human evaluators?
@@
To answer the above research questions, we conducted comprehensive experiments on two widely-used benchmark datasets, ReDial (Li et al., 2019) and OpenDialKG (Moon et al., 2019), involving four representative CRSs, including both neural-based models (BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), UniCRS (Wang et al., 2022b)) and a GPT-based model (CHATCRS (Wang et al., 2023a)). We adopted a modified user simulator based on Wang et al. (2023a) to generate multi-turn dialogue logs that realistically mimicked user interactions by providing vague requests and ensuring a dynamic conversation. The generated logs were then evaluated using various large language models (e.g., GPT-4o-mini, GLM-4-Air, the Llama-3 family, the Mistral family, and the Qwen-3 family) across the twelve human-centered factors. Then the Multi-Agent Debater based on GPT-4o-mini was introduced to aggregate factor-wise scores provided by GPT-4o-mini into a unified overall assessment score. Additionally, to validate the proposed framework, we conducted a human evaluation where ten independent annotators rated a total of 80 randomly selected dialogue logs from the test sets. Each evaluator scored the logs on the same twelve factors and provided an overall score, allowing for direct comparison with the LLM-based assessments.
From experimental results, we have the following key findings: (1) The key factors in the proposed framework show a degree of independence under human evaluation, but LLM-based evaluators tend to blur those boundaries to different degrees: some models introduce only moderate coupling, while others create much stronger inter-factor interdependencies. (2) The proposed framework (CoRE), with appropriate LLM serving as the evaluator, aligns well with human evaluation in most of the key factors influencing user experience in conversational recommendations. (3) CoRE, with appropriated LLM serving as the evaluator, aligns well with human evaluation in terms of CRS overall performance evaluation. (4) LLMs have strong potential of being used directly as CRS. Neural network-based CRS could exhibit deficiencies in some key user experience factors such as Semantic Relevance, Explainability, and Proactiveness, which negatively impact the overall user interaction experience.
@@ 2.1. Conversational Recommender Systems
Dialogue systems are designed to provide social support or functional services to human users through natural language interactions. Key areas in this field have primarily focused on dialogue context comprehension (Chen et al., 2022b) and response generation techniques (Roller et al., 2021). Recommender Systems aim to help users filter and select items that meet their information needs. To achieve this goal, traditional recommender systems generate recommendations based on implicit feedback from past user-item interactions (Schafer et al., 2007; Ekstrand et al., 2011; Koren, 2008; Bell and Koren, 2007; Koren et al., 2009; Covington et al., 2016). As the pace of modern life accelerates, traditional recommender systems face inherent limitations due to their static interaction paradigms and insufficient adaptability to users’ evolving preferences. In response, Conversational Recommender Systems (CRSs) have emerged as a promising research frontier, integrating intelligent dialogue agents with recommendation engines to enable dynamic, context-aware, and user-centric interactions (Ren et al., 2020). Taking a different approach to traditional RSs, CRSs aim to model the relationship between user preferences and candidate item representations to deliver accurate recommendations, while simultaneously generating coherent responses that are contextually aligned with the recommended items (Jannach et al., 2021; Gao et al., 2021; Zamani et al., 2023; Li et al., 2018; Liang et al., 2021). As CRSs interact with users through multi-turn dialogues, CRSs are also regarded as a specialized form of dialogue systems (Deng et al., 2025; Li et al., 2021; Ren et al., 2020).
@@
In our study, we focus on the evaluation of chit-chat based CRS, and we selected four representative CRSs for evaluation, in which three (BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019) and UniCRS (Wang et al., 2022b)) are based on neural networks, while one (CHATCRS (Wang et al., 2023a)) directly utilizes GPT as a CRS.
@@
We selected the above systems because many studies (Wang et al., 2023a, b; Zhu et al., 2024a, b; Bernard et al., 2025; Bernard and Balog, 2025) have evaluated them or used them as baselines, which helps to connect our work with existing evaluations of conversational recommender systems.
@@ 2.2. Evaluating Conversational Recommender Systems
Evaluation is a fundamental aspect of developing search and recommender systems, encompassing both user-centric assessments, which capture the effectiveness of the system from the user’s perspective, such as user satisfaction (Bernard et al., 2025); as well as system-centric metrics such as Precision, Recall, and inference time (Siro et al., 2023; Al-Maskari and Sanderson, 2010). The evaluation of CRS also encompasses both user-centric and system-centric approaches, and can be further categorized into three main scenarios:
  • Lab-based User studies examine perceived system quality, evaluating aspects like recommendation relevance, perceived transparency, and ease of use (Pu et al., 2011).
  • Online metrics, such as A/B testing, are commonly used in real-world settings to assess whether the system achieves its intended objectives, typically through metrics such as user engagement or conversion rates (Jannach and Jugovac, 2019).
@@
Due to the high cost of A/B testing and user studies, offline evaluation is commonly adopted in practice to assess CRSs. The most commonly used metrics for evaluating CRS include Recall, Precision, etc., for assessing recommendation task performance, and BLEU, ROUGE, Perplexity, Distinct-N, etc., for evaluating dialogue management task performance. Nevertheless, empirical studies indicate that these rule-based metrics do not align with human preferences (Chen et al., 2017; Reiter, 2018; Cremonesi et al., 2012). One key contribution of our study is the use of large language models (LLMs) as evaluators to enable user-centric offline evaluation.
@@
In the context of user-centered evaluation for CRSs, previous studies have argued that several factors may influence users’ assessment of their interactions with the system, from both the recommendation and dialogue perspectives . These factors include: (perceived) accuracy of recommendation (Contreras et al., 2021; Manzoor et al., 2024; Jin et al., 2021; Zhang et al., 2024c), diversity of recommended items (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024c), novelty of recommended items (Jannach and Bauer, 2020; Shani and Gunawardana, 2011; Jin et al., 2021; Zhang et al., 2024c), explainability of recommendation (Sakai, 2023; Jin et al., 2021; Zhang et al., 2024c, b), user intent understanding and response capability (e.g., maintaining a coherent dialogue flow, recovering from errors) (Sakai, 2023; Langevin et al., 2021; Jin et al., 2021; Zhang et al., 2024c), the quality of system-generated text (e.g., fluency, linguistic appropriateness and naturalness) (Jin et al., 2021; Jannach, 2022; Langevin et al., 2021), system informativeness for decision-making (e.g., information sufficiency, information accuracy) (Sakai, 2023; Jannach, 2022), proactiveness of the system (Kraus et al., 2020; Deng et al., 2025; Christakopoulou et al., 2016). Nevertheless, these studies did not propose any computable metric for evaluating CRS performance.
@@
Some prior work has explored both human-annotated and LLM-based approaches for evaluating CRSs from a user-centered perspective.  Manzoor and Jannach (2022) uses crowd-sourced pairwise comparisons to assess CRS response quality, and Yang et al. (2024) proposed a behavior alignment method that relies on human-annotated data. However, these methods rely on collecting human data, which is costly.  Wang et al. (2023a) utilized LLMs to assess whether the items recommended by CRS meet user needs, while  Huang et al. (2024) leveraged LLMs to evaluate the social attributes of CRS.  Zhang et al. (2024b) leveraged LLMs to assess the quality of explanation texts generated by CRSs. However, these studies did not propose a method for assessing the overall performance of CRS. To the best of our knowledge, our work is the first to propose a unified, LLM-based framework for holistic, user-centered evaluation of CRSs. Our work bridges the gap between user studies on CRSs and current evaluation practices in the field. We extracted twelve factors influencing users’ interaction experiences with conversational recommender systems from prior studies, used LLMs to assign evaluation scores to each factor, and then aggregated these into an overall score. By proposing a feasible evaluation framework, our study contributes to advancing CRS evaluation practices toward a more human-centered paradigm.
2.3. Large Language Models as Evaluators for Text Quality